OpenStudy (anonymous):

look at the following points. (4,0), (3,-1), (6,3), (2,-4) which are solutions to y=x-4 choose all the correct answers. -6,3 -4,0 -3,-1 -2, -4

OpenStudy (goformit100):

A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (amistre64):

do you know how to read the information that a point gives you?

OpenStudy (anonymous):

not really..

OpenStudy (amistre64):

points are defined for x and y values, they are read as: (x,y) take the first point for an example: (4,0) , this tells us that x=4 and y=0

OpenStudy (anonymous):

makes sense!

OpenStudy (amistre64):

in order to determine of it satisfies the given equation; just plug in the values to see if they produce a true result

OpenStudy (anonymous):

so would it be the last one?

OpenStudy (amistre64):

y = x-4 , try the point (4,0): x = 4, and y = 0 | | v v 0 = 4 - 4 does 0 = 4-4 ?

OpenStudy (anonymous):

yes!

OpenStudy (amistre64):

then (4,0) is a good point :) lets test the last one (2,-4), x=2 and y=-4 y = x-4 | | v v -4 = 2 - 4 does -4 = 2-4 ?

OpenStudy (anonymous):

no!

OpenStudy (amistre64):

then we would want to avoid that one since it doesnt fit .... test out the remiaing points in the same way to see if they fit or not

OpenStudy (anonymous):

x y 3,-1 y=x-4 3= -1 -4 false.

OpenStudy (anonymous):

that right?

OpenStudy (amistre64):

remember to keep an eye on where the x and y values get placed. I know its easy to mix them up.

OpenStudy (amistre64):

x y 3,-1 y = x-4 3= -1 -4 ^ ^ youve placed them in the wrong spots

OpenStudy (amistre64):

x y 3,-1 y = x -4 -1= 3 -4 ^ ^ this is how we want them to go; y for y and x for x

OpenStudy (amistre64):

since -1 = 3-4, this point is good

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

good luck :)